How many pounds of a 35% salt solution and how many pounds of a 14% salt solution should be combined so that 50 pounds of a 20% salt solution are obtained?
Answer by ptaylor(2198)


Let x=number of pounds of 35% solution
Then 50-x=number of pounds of 14% solution
Now we know that the pure salt in the 35% solution (.35x) plus the amount of pure salt in the 14% solution (.14(50-x)) has to equal the amount of pure salt in the final mixture (.20(50). So our equation to solve is:
.35x+.14(50-x)=.20(50) multiplying out:
.35x+7-.14x=10 subtract 7 from both sides
.35x+7-7-.14x=10-7 collect like terms
.21x=3
x=14.285 lb -------------number of pounds of 35% solution
50-x=50-14.285=35.715-------------number of pounds of 14% solution
Ck
substitute into original equation
.35(14.285)+.14(35.715)=.20(50)
5+5=10
10=10

Hope this helps-----ptaylor